In a non-leap year, the city of Montreal experienced 12 days less frost than the city of Quebec. In all, the two cities experienced 224 days of frost. How many frost-free days did the city of Montreal experience that year?
Answer by ikleyn(52781)

.

    Q - M = 12   days

    Q + M = 224  days


Add the equation and get  2Q = 12 + 224 = 236;  Q = 236/2 = 118.


Then from the second equation  M = 224 - 118 = 106.


ANSWER.  106 days for Montreal.

Solved.

------------

To see one million other similar solved problems, look into the lesson
    - Word problems that lead to a simple system of two equations in two unknowns
in this site.


/\/\/\/\/\/\/

Notice that the fact whether a year is a leap year or not a leap year is   I R R E L E V A N T   to this problem.